OpenStudy (anonymous):

2 cosx = x + (1/x) then value of x^6 - (1/x^6) is?

OpenStudy (experimentx):

find the value of x^3 + 1/x^3 and x^3 - 1/x^3

OpenStudy (anonymous):

\[z+\frac{1}{z}=2\cos x\]\[z^6-\frac{1}{z^6}=?\]

OpenStudy (experimentx):

square it ... try some manipulation ... ( x - (1/x))^2 +2 = ? a^3 + b^3 = ? a^3 - b^3 = ? make use of these basic simple formulas.

OpenStudy (anonymous):

or one can write\[z=e^{ix}\]and find the value of\[z^6-z^{-6}\]

OpenStudy (experimentx):

nice observation ... didn't think of that.

OpenStudy (anonymous):

@mukushla i cant see ur...Witings....

OpenStudy (anonymous):

reload the page

OpenStudy (anonymous):

lol....i mean....Font is Small

OpenStudy (anonymous):

lol\[\large z=e^{ix}\]find the value of\[\large z^6-z^{-6}\]

OpenStudy (anonymous):

e^ix - 1/e^ix

OpenStudy (anonymous):

\[\large z^6-z^{-6}=e^{6xi}-e^{-6xi}=\cos 6x+i\sin 6x-(\cos 6x-i\sin 6x)\]is this right? i have some doubts on it

OpenStudy (anonymous):

Yes...u r correct

OpenStudy (anonymous):

so it gives answer as \[2i \sin 6x\]